OpenStudy (anonymous):

What values for q (0 £ q < 2p) satisfy the equation? 2√2sinq + 2 = 0

OpenStudy (anonymous):

\[2\sqrt{2}\sin(q)+2=0\]

OpenStudy (anonymous):

\[2\sqrt{2}\sin(q)=-2\] \[\sqrt{2}\sin(q)=-1\] \[\sin(q)=-\frac{1}{\sqrt{2}}=-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

now what you need is a number \(q\) with \[ \sin(q)=-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

find it on the attached cheat sheet that is, find where the second coordinate is \(-\frac{\sqrt{2}}{2}\) the angle will be your answer

OpenStudy (anonymous):

that wasn't one of my answers.